OpenStudy (anonymous):

6/(y-5) - (Y+5)/(Y^2-25)

OpenStudy (anonymous):

its not a full question

OpenStudy (anonymous):

Factor.

OpenStudy (anonymous):

where is factor sign and there shud be an equation sign

OpenStudy (anonymous):

The expresion above is equivalent to: (Thats the question)

OpenStudy (tkhunny):

Two fractions. Add them. \(y^{2}-25 = (y+5)(y-5)\)

OpenStudy (anonymous):

But after that what do you do next? The denominators are still different

OpenStudy (tkhunny):

Add fractions. You learned how to do that a long time ago. It is time to prove it. \(\dfrac{1}{4} + \dfrac{1}{8} = \dfrac{1}{4}\cdot\dfrac{2}{2} + \dfrac{1}{8} = \dfrac{2}{8} + \dfrac{1}{8} = \dfrac{3}{8}\) That should look familiar. That is what you need to do.